The difference between the sum of even numbers and the sum of odd numbers between 10 and 20 is
Answer by ikleyn(53765)

.
The difference between the sum of even numbers and the sum of odd numbers between 10 and 20 is
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

  10 +   12 +   14 +   16 +   18  +  20
 -
  11 +   13 +   15 +   17 +   19 
--------------------------------------

 (-1) + (-1) + (-1) + (-1) + (-1) +  20  = -5 + 20 = 15.


Grouping !